OpenStudy (anonymous):

Find the Derivative P(t) = 2*e^(-2e)^(2t) P prime (t) =

OpenStudy (anonymous):

By using the chain rule we get\[P'(t)=2e^{(-2e)^{2t}}*-2e^{2t}*2\] It's a tedious one, but if you take it slow and break it down and it's not too bad. Chain rules within chain rules are always annoying

OpenStudy (anonymous):

yeah, it marked it as wrong ^^^^^

OpenStudy (anonymous):

Interesting.

zepdrix (zepdrix):

If the exponent is `really` (-2t)^(2t), then this one is going to be be a lil more frustrating since the -2 is included lol

OpenStudy (anonymous):

yeah ^

OpenStudy (anonymous):

would we make a substitution? \[u=-2e ^{2t}\]

OpenStudy (anonymous):

See that's different than what you initially wrote :P

zepdrix (zepdrix):

\[\large\rm y=a^x\]Recall that for differentiating exponentials:\[\large\rm y'=a^x(\ln a)\]So then,\[\large\rm \frac{d}{dt}(-2e)^{2t}=(-2e)^{2t}\cdot \ln(-2e)\cdot (2)\]Oh but then log of -2e isn't defined... Yah having a negative base is uhh... no bueno :U hmm

OpenStudy (anonymous):

yeaah, see i tried wolfphram alpha and it gave me an imaginary number???

zepdrix (zepdrix):

Do you have a picture of the problem or something?

zepdrix (zepdrix):

Ugh :( Ya that's not what you wrote out on top -_-

OpenStudy (anonymous):

sorry i tried my best

zepdrix (zepdrix):

Then Chris's derivative is correct :) We just gotta be careful with the brackets.

zepdrix (zepdrix):

P(t)=2*e^(-2e^(2t)) P'(t)=2e(-2e^(2t))*(-2)e^(2t)*2 Try that maybe? :d Do you only get a limited number of "guesses"?

OpenStudy (anonymous):

I'm on my last one, but oh well only got 3 mins left

zepdrix (zepdrix):

oh XD

zepdrix (zepdrix):

oops i made a typo :O hopefully you caught that. umm the first part of it

zepdrix (zepdrix):

P'(t)=2e^(-2e^(2t))*(-2)e^(2t)*2 there we go, forgot the carot on the first exponential >.<

OpenStudy (anonymous):

carot hahaaha

zepdrix (zepdrix):

caret* :3 i dunno, whatever this thing is called "^" hehe

OpenStudy (anonymous):

oh haha it sorta looks like a carrot anyway

OpenStudy (anonymous):

ouch lol was a second late, when i hit submit it turned to 12 am :(

OpenStudy (anonymous):

automatically took me home

zepdrix (zepdrix):

aw :P does it show solution?

zepdrix (zepdrix):

lame t.t

OpenStudy (anonymous):

it doesn't :(

OpenStudy (anonymous):

im gonna get on it on office hours, thank guys!